OpenStudy (angelina__nievse):

Simplify 11 exponent -4/ 11 exponent 8 please help will give medal

OpenStudy (sachintha):

\(\huge \frac{a^b} {a^c}=a^{b-c}\) So, \(\Large\frac{11^{-4}} {11^8}=11^{-4-8}\)